OpenStudy (anonymous):

A bag contains 4 nickels, 8 dimes, and 2 quarters. If you draw 3 coins at random from the bag, without replacement, what is the probability that you will get a nickel, a quarter, and a nickel, in that order?

OpenStudy (abbles):

the probability of getting a nickel first is 4/14 which simplifies to 2/7. Then, the probability of getting a quarter is 2/13. The probability of drawing a nickel the third time is 3/12 which simplifies to 1/4. Now, you need to multiply all of the fractions together. 2/7 x 2/13 x 1/4 = 1/91 which is your answer.

OpenStudy (anonymous):

\[\frac{4}{14}*\frac{8}{13}*\frac{2}{12}=\frac{8}{273} \]

OpenStudy (abbles):

@robtobey the third fraction is incorrect... there are three nickels to be drawn, not two. Also, on the second fraction there are only 2 quarters, not 8. Otherwise, it looks good.